# A Short History of (Orbital) Decay: Roman’s Prospects for Detecting Dying Planets

### v1

@@ 4.3 Alternative Parameterization
The exact form of $Q_{*}^{{}^{\prime}}$ is not well-constrained. Millholland et al. (2025) argued using the known population of short period planets that
@@
with $-4.33\lesssim\alpha\lesssim-2$ and $10^{5.5}\lesssim Q_{0}\lesssim 10^{7}$, suggesting less efficient tidal dissipation at short periods, consistent with Zanazzi et al. (2024). We adopted $\alpha=-3$ and $Q_{0}=10^{6}$ and then re-estimate the SNR distribution, as shown in Fig. 7.
Figure 7: The histogram of orbital decay SNRs $>10^{-3}$ using the $Q_{*}^{{}^{\prime}}$ parameterization from Millholland et al. (2025).
@@ 4.4 Caveats
Planet occurrence rates and properties in the Roman field might differ greatly from what we observe in the Solar neighborhood. The yield of orbital decay is likely to be especially sensitive to how many transiting planets Roman observes orbiting subgiant stars. Roman will set lower limits on $Q_{*}^{{}^{\prime}}$ for an unprecedentedly large sample of transiting planet hosts, improving empirical constraints on $Q_{*}^{{}^{\prime}}$.
For a variety of factors, our yield estimate is likely overly conservative. First, the Wilson et al. (2023) catalog used occurrence rates with only upper limits for much of the parameter space where $P\lesssim 1$ day. We expect that the catalog underestimates the number of planets that will be detected in the regime of interest for orbital decay. Additionally, the catalog does not simulate planets around giants, which Roman is expected to detect.
Orbital decay is not the only phenomenon that causes deviations from a linear ephemeris model. Precession, dynamical perturbations from other bodies, and acceleration of the system along our line of sight can lead to a non-zero period derivative. Since dynamical interactions with other planets would be periodic, line-of-sight acceleration and apsidal precession are the most nefarious false positive scenarios. It often requires some detective work to decipher the physical origin of such a signal (e.g., Bouma et al., 2020; Ivshina & Winn, 2022). Radial velocity measurements will be difficult to acquire for Roman stars so excluding line-of-sight acceleration might be difficult. One can exclude apsidal precession by timing secondary eclipses (Yee et al., 2020) or by inferring precession would require unphysically large planet Love numbers (Vissapragada et al., 2022). The latter of these approaches is likely to be more feasible with Roman stars.
@@ 5 Conclusions
We expect that Roman will be a powerful laboratory for probing orbital decay, a key mechanism for sculpting planetary demographics in the small-period regime. We expect $5-10$ examples of orbital decay to be detectable over the entire duration of the GBTDS. If these detections do come to fruition, they would expand the number of firm orbital decay detections by a factor of 2 or more.
@@ 6 Acknowledgments
K.C. would like to thank Luke Bouma, Sarah Millholland, and Christopher Kochanek for useful comments and feedback. K.C., B.S.G., and R.F.W. were supported by NASA Grant 80NSSC24K0917. B.S.G. was additional supported by the Thomas Jefferson Chair Endowment for Discovery and Space Exploration, and K.C. was additionally supported by the Ohio State Distinguished University Fellowship. The material is based upon work supported by NASA under award number 80GSFC24M0006.

### v2

@@ 4.3 Alternative Parameterization
The exact form of $Q_{*}^{{}^{\prime}}$ is not well-constrained. millholland_empirical_2025 argued using the known population of short period planets that
@@
with $4.33\lesssim\alpha\lesssim-2$ and $10^{5.5}\lesssim Q_{0}\lesssim 10^{7}$, suggesting less efficient tidal dissipation at short periods, consistent with zanazzi_damping_2024. We adopted $\alpha=-3$ and $Q_{0}=10^{6}$ and then re-estimate the SNR distribution, as shown in Fig. 7.
Figure 7: The histogram of orbital decay SNRs $10^{-3}$ using the $Q_{*}^{{}^{\prime}}$ parameterization from millholland_empirical_2025.
@@ 4.4 Caveats
Planet occurrence rates and properties in the Roman field might differ greatly from what we observe in the Solar neighborhood. Three of the four high-significance systems shown in Table 1 are among the $\sim 200$ catalog systems with log(gg)<3.8\,<3.8, suggesting the yield of orbital decay is especially sensitive to how many transiting planets Roman observes orbiting subgiant stars. Roman will set lower limits on $Q_{*}^{{}^{\prime}}$ for an unprecedentedly large sample of transiting planet hosts, improving empirical constraints on $Q_{*}^{{}^{\prime}}$.
For a variety of factors, our yield estimate is likely overly conservative. First, the wilson_transiting_2023 catalog used occurrence rates with only upper limits for much of the parameter space where $P\lesssim 1$ day. We expect that the catalog underestimates the number of planets that will be detected in the regime of interest for orbital decay. Additionally, the catalog does not simulate planets around giants, which Roman is expected to detect.
Orbital decay is not the only phenomenon that causes deviations from a linear ephemeris model. Precession, dynamical perturbations from other bodies, and acceleration of the system along our line of sight can lead to a non-zero period derivative. It often requires some detective work to decipher the physical origin of an inferred period derivative (e.g., bouma_wasp-4_2020; ivshina_tess_2022). Radial velocity measurements will be difficult to acquire for Roman stars so excluding line-of-sight acceleration directly is generally not feasible. The number of detectable period derivatives caused by line-of-sight acceleration depends on the number of widely-orbiting companions to planet-hosting stars and on their eccentricity distribution. Line-of-sight acceleration should produce systems whose periods are inferred to be increasing as well as decreasing so observing a substantial excess of inferred negative period derivatives would help establish orbital decay at a population-level. One can exclude apsidal precession by timing secondary eclipses (yee_orbit_2020) or by inferring precession would require unphysically large planet Love numbers (vissapragada_possible_2022). Apsidal precession causes a sinusoidal, rather than quadratic, timing deviations so long-term transit time modeling might rule out precession as well. It might be difficult to establish orbital decay as the most likely cause of an inferred period derivative for any individual system.
@@ 5 Conclusions
We expect that Roman will be a powerful laboratory for probing orbital decay, a key mechanism for sculpting planetary demographics in the small-period regime. We expect 5105-10 examples of orbital decay to be detectable over the entire duration of the GBTDS. If these detections do come to fruition, they would expand the number of firm orbital decay detections by a factor of 2 or more.
@@ 6 Acknowledgments
We would like to thank our anonymous referee, whose suggestions substantially improved the quality of this paper. K.C. would like to thank Luke Bouma, Sarah Millholland, and Christopher Kochanek for useful comments and feedback. K.C., B.S.G., and R.F.W. were supported by NASA Grant 80NSSC24K0917. B.S.G. was additionally supported by the Thomas Jefferson Chair Endowment for Discovery and Space Exploration, and K.C. was additionally supported by the Ohio State Distinguished University Fellowship. The material is based upon work supported by NASA under award number 80GSFC24M0006.